OpenStudy (anonymous):

how do i check t/4-t/5=1 (t=20)

OpenStudy (akshay_budhkar):

jus substitute t=20... you get 5-4=1=lhs=rhs so you are right:)

OpenStudy (anonymous):

20/4 = 5 20/5 = 4 5 -4 = 1